OpenStudy (jb1625):

What property is shown in the equation? 5 ∙ (0.6ab) = 3ab A. Associative Property of Multiplication B. Commutative Property of Multiplication C. Identity Property of Multiplication D. Reciprocal Property of Multiplication

OpenStudy (jango_in_dtown):

5*(0.6ab)=(5*0.6)ab=3ab, associative property of multiplication

OpenStudy (jango_in_dtown):

@JB1625

OpenStudy (jb1625):

Ok thank you

OpenStudy (jb1625):

@jango_IN_DTOWN